Show that the tangents from the point with position vector −2i − 3j to the ellipse 4x2 + 9y2 = 36 are perpendicular.


1
Expert's answer
2021-08-16T15:17:46-0400
"4x^2+9y^2=36"

Differrentiate borh sides with respect to "x"


"8x+18yy'=0"

"y'=-\dfrac{4x}{9y}"

"slope=m=-\dfrac{4x_0}{9y_0}"

The equation of the tangent to the ellipse is


"y-y_0=m(x-x_0)"

"y=mx-mx_0+y_0"

"x_0=-\dfrac{9y_0}{4}m"

"y_0-mx_0=\dfrac{y_0(4+9m^2)}{4}"

"4x_0^2+9y_0^2=36"

"\dfrac{81y_0^2}{4}m^2+9y_0^2=36"

"y_0^2(9m^2+4)=16"

"\dfrac{y_0}{4}=\pm\dfrac{1}{\sqrt{9m^2+4}}"

"y=mx\pm\sqrt{9m^2+4}"


Point "(-2, -3)"


"-3=m(-2)\pm\sqrt{9m^2+4}"

"2m-3=\pm\sqrt{9m^2+4}"

"4m^2-12m+9=9m^2+4"

"5m^2+12m-5=0"

"m_1m_2=\dfrac{-5}{5}=-1"

Therefore the tangents from the point with position vector "-2i-3j" to the ellipse "4x^2+9y^2=36" are perpendicular.